# SQ-LLaVA: Self-Questioning for Large Vision-Language Assistant

### v1

@@ Abstract
Recent advancements in the vision-language model have shown notable generalization in vision-language tasks after visual instruction tuning. However, bridging the gap between the pre-trained vision encoder and the large language models becomes the whole network’s bottleneck. To improve cross-modality alignment, existing works usually consider more visual instruction data covering a broader range of vision tasks to fine-tune the model for question-answering, which are costly to obtain. However, the image contains rich contextual information that has been largely under-explored. This paper first attempts to harness this overlooked context within visual instruction data, training the model to self-supervised ‘learning’ how to ask high-quality questions. In this way, we introduce a novel framework named SQ-LLaVA: Self-Questioning for Large Vision-Language Assistant. SQ-LLaVA exhibits proficiency in generating flexible and meaningful image-related questions while analyzing the visual clue and prior language knowledge, signifying an advanced level of generalized visual understanding. Moreover, fine-tuning SQ-LLaVA on higher-quality instruction data shows a consistent performance improvement compared with traditional visual-instruction tuning methods. This improvement highlights the efficacy of self-questioning techniques in achieving a deeper and more nuanced comprehension of visual content across various contexts. Our implementation is available at https://github.com/heliossun/SQ-LLaVA.
@@ 1 Introduction
The recently emerging large vision-language methods, such as large language-and-vision assistant (LLaVA) and its variants Liu et al. (2023a, b); Chen et al. (2023); Bai et al. (2023); Zhang et al. (2024), fine-tune large language models (LLM) Zheng et al. (2023); Zhang et al. (2024); Bai et al. (2023) on visual instruction data to realize diverse open-world multimodal understanding, demonstrating a surprising efficacy of visual instruction tuning – the LLM learns to perform complex vision tasks simply by conditioning on a prompt containing image and text clues. Existing visual instruction data is mainly built upon conversations (e.g., ChatGPT/GPT4-V 202 (2023)), each consists of an image and multiple question-answer pairs. Building high-quality visual instruction data usually requires images and texts from different tasks to generate diverse questions, such as “Please provide the bounding box coordinate of the region this sentence describes A dead leaf on the ground.” for the object detection task. Empirically, by increasing the diversity of questions, LLaVA has achieved better performance on GQA and VizWiz tasks (26% and 45% over previous state-of-the-art methods Dai et al. (2023)). This evidence strongly suggests the advantage of training models on a broad spectrum and diverse array of tasks for enriching general vision-language understanding.
(a) Training strategy comparison.
(b) Performance comparison.
@@
Figure 1: A high-level comparison between visual instruction tuning and visual self-questioning (ours) for vision-language assistant. SQ-LLaVA achieves state-of-the-art performance on 9 out of 10 tasks compared with other open-ended models.
@@
LLaVA Liu et al. (2023b) model family usually consists of a pre-trained vision encoder (e.g., CLIP-ViT Radford et al. (2021)), a large generative language model (e.g., Vicuna Zheng et al. (2023), LLaMA Zhang et al. (2024), and QWen Bai et al. (2023)), and a vision-to-language projector implemented by a few linear layers. However, the modality gap between the pre-trained vision encoder and the language model restricts both sides’ generalization ability and feature representation. To overcome this challenge, various techniques have been proposed to align the vision and language domains, which could be roughly categorized into three groups: 1) build a more robust image feature extractor  Zhang et al. (2024); Zhu et al. (2024); Dai et al. (2023), 2) collect more high-quality training data Liu et al. (2023b); Chen et al. (2023); Zhu et al. (2024); Dai et al. (2023), and 3) fully fine-tune the vision and language models simultaneously during the pre-training stage Chen et al. (2023). While these methods have shown good progress in mitigating the domain gap, they inevitably bring higher computational costs and more expensive data collection and may also require sophisticated manual designs and heavy annotating efforts. Moreover, the image encompasses a wealth of information, including color, context, and the relationships between objects, etc., but most visual instruction datasets capture only a fraction. We posit that leveraging such under-explored knowledge could significantly aid in multi-modal understanding. In this study, we propose a new visual self-questioning approach by training the model to ask questions and discover vision clues without collecting extra data from other sources.
Unlike the existing visual instruction tuning methods, which treat the answer prediction as the only training target, the proposed visual self-questioning aims to mine the informative context of questions. For example, as shown in Fig. 0(a), the questions: “What is the person wearing on there feat?” and “How many bears are there?” has the answer “Socks” and “1”. Obviously, the question contains more image-related information missed by the answer, so only training the model to predict the answer under this situation has no sufficient guidance to optimize the general vision-language understanding. To this end, due to the diverse semantic information, questions within current visual instruction data can be utilized to fine-tune the instructional LLM. Therefore, we introduce a coupled visual self-questioning task besides learning the question-answering tasks.
@@
This work proposes a novel visual self-questioning and instruction tuning model, SQ-LLaVA, utilizing the questions within the instruction data as another learnable resource to train the instructional LLMs (Fig. 0(a)). To jointly align the vision and language domains efficiently, we apply LoRAs Hu et al. (2022) to the vision encoder and the LLM and optimize them together. SQ-LLaVA proposes a prototype extractor to enhance visual representation since the learned clusters with meaningful semantic information can improve vision language alignment. Through extensive experiments (Fig. 0(b)), we show that SQ-LLaVA can further improve general vision understanding than existing visual instruction tuning methods. We summarize the contributions as follows.
@@
  • We propose a new training technique called visual self-questioning for vision-language assistant (SQ-LLaVA) by taking advantage of informative question context in instruction data. This new SQ learning task encourages instructional LLM to capture the relationship between images and questions, leading to better vision-language alignment without collecting new data.
@@
  • We develop a lightweight architecture of the vision-language model for SQ-LLaVA. Specifically, the prototype extractor effectively enhances the vision embedding with meaningful prototype information, and the ViT-LoRA and LLM-LoRA efficiently align vision and language domains during training.
@@
  • Extensive experimental results show that our proposed method leads to better performance in several areas. In traditional Visual Question-answering tasks, we achieve $2.5\%$ improvement over the LLaVA-v1.5. For visual instruction tasks, our improvement is $7.2\%$. Moreover, we improve the performance by $2\%$ in cross-domain image captioning.
@@ 2.1 Instruction Tuning
Instruction tuning emerged as a pivotal methodology within the realm of natural language processing (NLP), facilitating Large Language Models (LLMs) such as GPT-3 Brown et al. (2020), PaLM Chowdhery et al. (2022), and LLaMA Touvron et al. (2023a) to interpret and execute human language instructions across a spectrum of NLP tasks. This approach diverges from traditional fine-tuning mechanisms by incorporating a specialized data structure, termed instruction-following data Wei et al. (2022), which is instrumental in the fine-tuning process of LLMs. Predominantly, there are two principal methodologies concerning instruction tuning. Early scholarly contributions delineated the initial methodology, closed-domain instruction tuning Wei et al. (2022); Sanh et al. (2022); Xu et al. (2024). These studies engaged LLMs with a comprehensive assortment of publicly accessible datasets and subsequently assessed their performance across diverse NLP tasks Triantafillou et al. (2019). The empirical evidence from these inquiries consistently indicated that integrating varied NLP task instructions significantly augments the LLMs’ efficacy in navigating novel tasks. Nonetheless, LLMs calibrated with such closed-form instructions exhibited limitations in real-world user scenarios, prompting the development of an alternative approach. To address these constraints, the concept of open-domain instruction tuning Ouyang et al. (2022a); Wang et al. (2022a) was conceived. OpenAI pioneered this approach by employing human annotators to compile a corpus of real-world question-answer datasets. These datasets were the foundation for training a reward model through reinforcement learning methodologies. The trained reward model then functioned as a supervisory mechanism for further training instruction-oriented language models, such as InstructGPT Ouyang et al. (2022b) and Vicuna Zheng et al. (2023). This innovation marked a significant advancement in the field, aiming to bridge the gap between LLM performance and real-world applicability by leveraging instruction data derived from authentic user interactions.
@@ 2.2 Large Vision Language Models
As the field of LLMs and instruction tuning undergoes rapid advancements, the academic research community is progressively focusing on integrating visual information into LLM frameworks to facilitate visual instruction tuning. This emerging area of research has witnessed the proposal of various methodologies based on the foundational work of CLIP Radford et al. (2021) and diverse LLM Zheng et al. (2023); Touvron et al. (2023b); Bai et al. (2023) architectures. Notably, LLaVA Liu et al. (2023a) pioneered the integration of an LLM with a CLIP vision encoder to construct a vision-language model, demonstrating remarkable capabilities in image-text dialogue tasks through strategies of pre-training alignment and targeted instruction tuning. Subsequent investigations have sought to refine visual instruction tuning by enhancing the quality and variety of the datasets used during the pre-training and fine-tuning phases. Building upon these advancements, recent studies, including LLaVA-v1.5 Liu et al. (2023b) and ShareGPT4V Chen et al. (2023), have achieved notable success in general vision-language comprehension, showcasing their ability to undertake complex question-answering tasks. This progression underscores the importance of sophisticated data handling and model-tuning strategies in developing effective vision-language models.
@@ 3.1 Architecture Overview
The proposed SQ-LLaVA model (see Fig. 2) consists of four main components: 1) A pre-trained vision encoder CLIP-ViT Radford et al. (2021) that extracts a sequence embedding of image tokens $Z_{v}$ for an input image $X_{v}$; 2) A prototype extractor $\phi(\cdot)$ learning visual clusters to enhance the original image tokens; 3) A trainable projection block $W(\cdot)$ with two linear layers to map the enhanced image tokens to the language domain tokens $H_{v}$, handling the dimension misalignment between the vision and language domain; and 4) Our LLM backbone $f(\cdot)$ implemented by the pre-trained Vicuna Zheng et al. (2023) to predict the next token upon the previous embedding sequence. Given the input question $X_{q}$ and answer $X_{a}$, a word embedding matrix is used to map them to contextual embeddings $H_{q}$ and $H_{a}$, and the distribution over $H_{a}^{(i+1)}$ can be obtained following the auto-regressive model as:
@@
where $\theta$ represents all the trainable parameters in our model, $\sigma(\cdot)$ is a softmax function, and $f(\cdot)$ outputs the last token embedding of the whole sequence. We denote $p_{\theta}$ as the prediction probability for the anticipated answer token $H_{a}^{(i+1)}$ at the position $i+1$, conditioning on the input image token embedding $H_{v}$, the question token embedding $H_{q}$, and the previous answer token embeddings $H_{a}^{(1:i)}$. As shown in Eq. (1), the proposed SQ-LLaVA applies the language model $f(\cdot)$ to model $p_{\theta}$ given by $H_{v}$, $H_{q}$, and $H_{a}^{(1:i)}$. Existing visual instruction tuning methods Liu et al. (2023b); Chen et al. (2023) are only able to predict $H_{a}$, yet cannot fully exploit the rich semantic clues within $H_{q}$. In this study, we propose visual self-questioning instructions to guide LLMs in capturing the visual content within questions.
@@ 3.2 Visual Self-questioning Instruction
In broad real-world scenarios, proactively asking a question requires more understanding and background knowledge than answering Tofade et al. (2013). Similarly, this work proposes visual self-questioning to encourage the LLM to discover deeper vision-language alignment and improve the overall instruction-following performance. Particularly, SQ-LLaVA treats questioning as a new learning objective beyond answering, which, to the best of our knowledge, is the first practice in the field of visual instruction tuning. While the current vision-language model can ask questions 202 (2023), such skill is still learned from question-answering through instruction tuning. However, our proposed method shows that the decoder-based LLM has the potential to learn more skills, such as how to ask questions spontaneously when a unique instruction token is given (e.g., we define [vusr] in our work). Furthermore, visual self-questioning can potentially improve general vision-language understanding. To be specific, as shown in Fig. 1, there are a certain amount of questions containing more meaningful image-related information than answers in the existing visual instruction data Liu et al. (2023b); Chen et al. (2023). Thus, we hypothesize that the vision-language understanding can be improved once the LLM learns how to predict informative questions about a given image.
@@
$X_{c}^{j}=\left\{\begin{matrix}(\texttt{[usr]},X_{q}^{(j)},\texttt{[aswr]},X_{a}^{(j)})&~{}~{}j=1\mathrm{~{}or~{}}j>1,R<\delta\\ ~{}(\texttt{[{\color[rgb]{0.5625,0.66796875,0.86328125}vusr}]},X_{q}^{(j)},\texttt{[aswr]},X_{a}^{(j)})&j>1,R>\delta\end{matrix}\right.,$ (2)
@@
SQ-LLaVA performs zero-shot questioning without any in-context knowledge or human language instruction since the only instructional prompt is a unique token [vusr]. After visual self-questioning on instruction data with various question formats $X_{q}$, the questions sampled by SQ-LLaVA are more diversified than GPT4-V (as shown in Fig. 4), since the LLM has learned alignment between image and questions, which is different from previous works 202 (2023); Wang et al. (2022b). Specifically, previous general-purpose vision language models such as GPT4-V 202 (2023) can generate questions based on a given image, but it requires explicit language instruction such as “Ask complex questions that are relevant to the content in the image”. Accordingly, the quality of generated questions highly relies on prompt engineering. Also, self-instruct Wang et al. (2022b) utilizes in-context learning to prompt the LLM to ask specific questions.

### v2

@@ Abstract
Recent advances in vision-language models have shown notable generalization in broad tasks through visual instruction tuning. However, bridging the gap between the pre-trained vision encoder and the large language models (LLMs) becomes the whole network’s bottleneck. To improve cross-modality alignment, existing works usually consider more visual instruction data covering a broader range of vision tasks to fine-tune the model for question-answering, which, however, is costly to obtain and has not thoroughly explored the rich contextual information contained in images. This paper first attempts to harness the overlooked context within visual instruction data, training the model to self-supervised “learning” how to ask high-quality questions. In this way, we introduce a novel framework named SQ-LLaVA: Self-Questioning for Large Vision-Language Assistant. SQ-LLaVA exhibits proficiency in generating flexible and meaningful image-related questions while analyzing the visual clue and prior language knowledge, signifying an advanced level of generalized visual understanding. Moreover, fine-tuning SQ-LLaVA on higher-quality instruction data shows a performance improvement compared with traditional visual-instruction tuning methods. This improvement highlights the efficacy of self-questioning techniques in achieving a deeper and more nuanced comprehension of visual content across various contexts. Our code is available at https://github.com/heliossun/SQ-LLaVA.
Keywords:
Vision-Language Understanding Multi-modal LLM Instruction Tuning
@@ 1 Introduction
The recently emerging large vision-language methods, such as large language-and-vision assistant (LLaVA) and its variants [22, 21, 5, 3, 54], fine-tune large language models (LLM) [57, 54, 3] on visual instruction data [56, 58, 55, 18] to realize diverse open-world multimodal understanding, demonstrating a surprising efficacy of visual instruction tuning – the LLM learns to perform complex vision tasks by conditioning on a prompt containing image and text clues. Existing visual instruction data is mainly built upon conversations (e.g., ChatGPT/GPT4-V [1]), consisting of images and multiple question-answer pairs. Building high-quality visual instruction data usually requires images and texts from different tasks to generate diverse questions, such as “Please provide the bounding box coordinate of the region this sentence describes A dead leaf on the ground” for the object detection task. Empirically, by increasing the diversity of questions, LLaVA has achieved better performance on GQA and VizWiz tasks (26% and 45% over previous state-of-the-art methods [8]). This evidence strongly suggests the advantage of training models on a broad spectrum and diverse array of tasks for enriching general vision-language understanding.
(a) Training strategy comparison
(b) Performance comparison
@@
Figure 1: (a) Comparision between visual instruction tuning and visual self-questioning (ours) for vision-language assistant. (b) The proposed SQ-LLaVA achieves state-of-the-art performance on 9 out of 10 tasks compared with other open-ended models.
@@
LLaVA [21] model family usually consists of a pre-trained vision encoder (e.g., CLIP-ViT [33]), a large generative language model (LLMs like Vicuna [57], LLaMA [54], QWen [3], etc.), and a vision-to-language projector implemented by a few linear layers. However, the modality gap between the pre-trained vision encoder and the language model restricts both sides’ generalization ability and feature representation. To overcome this challenge, various techniques have been proposed to align the vision and language domains, which could be roughly categorized into three groups: 1) build a more robust image feature extractor  [54, 58, 8], 2) collect more high-quality training data [21, 5, 58, 8], and 3) fully fine-tune the vision and language models simultaneously during the pre-training stage [5]. While these methods have shown good progress in mitigating the domain gap, they inevitably bring higher computational costs and more expensive data collection and may also require sophisticated manual designs as well as heavy annotating efforts. Moreover, images generally encompass rich information, including color, context, and the relationships between objects, but most existing visual instruction datasets capture only a fraction. We posit that leveraging such under-explored knowledge could significantly aid in vision-language understanding. In this study, we propose a new visual self-questioning approach (Fig. 1(a)) by training the LLM to ask questions and discover vision clues without collecting extra data from other sources.
Figure 2: Left: Questions spread more samples around a higher mean CLIPScore than answers. Right: Example of highly image-relevant questions within the visual instructional dataset for training a Vision-Language assistant.
Unlike existing visual instruction tuning methods that focus solely on answer prediction, the proposed visual self-questioning aims to extract relevant question context. As illustrated in Fig. 2 right, the questions could contain more image-related information than the answers, inspiring us that aligning images and related questions may further improve the model’s vision-language understanding capacity. Quantitatively, we compute the CLIPScore [33] (a higher value means better visual-text relevance) for all the image-question/-answer pairs on two visual instruction datasets (LLaVA-instruct [21] and ShareGPT4V-instruct [5]). By comparison, the questions’ mean CLIPScore $\mu_{q}$=$0.184$ is larger than the answers’ $\mu_{a}$=$0.183$ on LLaVA-instruct; $\mu_{q}$=$0.186$ is larger than the answers’ $\mu_{a}$=$0.184$ on ShareGPT4V-instruct. Fig. 2 left shows the CLIPScore distribution, indicating questions have a similar and even better visual relevance than the answers. Therefore, questions in current visual instruction data can be used to fine-tune instructional LLMs owing to their diverse semantic information.
@@
This work introduces a self-questioning LLaVA, namely SQ-LLaVA, to fully utilize questions within the instruction data as an additional learning resource for training instructional LLMs and empowering the model’s curiosity (questioning ability). To efficiently align the vision and language domains, we apply LoRAs[12] to optimize both the vision encoder and the instructional LLM within the SQ-LLaVA. Plus, we develop a prototype extractor to enhance visual representation by leveraging learned clusters with meaningful semantic information to improve vision-language alignment further. Extensive experiments demonstrate that SQ-LLaVA surpasses existing visual instruction tuning methods in general vision understanding (see Fig. 1(b)). We summarize the contributions as follows.
@@
  • We propose a novel training technique, visual self-questioning for vision-language assistants (SQ-LLaVA), by leveraging highly relevant question contexts in instruction data. This SQ learning task promotes instructional LLMs to understand the relationship between images and questions, enhancing vision-language alignment without needing new data collection.
@@
  • We design and develop a lightweight tuning architecture for SQ-LLaVA, consisting of ViT-LoRA, LLM-LoRA, and a prototype extractor. The prototype extractor enhances vision embeddings, while ViT-LoRA and LLM-LoRA efficiently align vision and language domains during training.
@@
  • Extensive experimental results show that the proposed SQ-LLaVA leads to better performance in several tasks, including visual question-answering, visual instruction benchmarks, and zero-shot image captioning.
@@ 2.1 Instruction Tuning
Instruction tuning emerged as a pivotal methodology within the realm of natural language processing (NLP), facilitating Large Language Models (LLMs) such as GPT-3 [4], PaLM [7], and LLaMA [42] to interpret and execute human language instructions across a spectrum of NLP tasks. This approach diverges from traditional fine-tuning mechanisms by incorporating a specialized data structure, termed instruction-following data [50], which is instrumental in the fine-tuning process of LLMs. Generally, there are two main categories regarding instruction tuning methods – 1) closed-domain and 2) open-domain. The closed-domain instruction tuning [50, 35, 51] studies engaged LLMs with a comprehensive assortment of publicly accessible datasets and subsequently assessed their performance across diverse NLP tasks [44]. The empirical evidence from these inquiries consistently indicated that integrating varied NLP task instructions significantly augments the LLMs’ efficacy in navigating novel tasks. Nonetheless, LLMs calibrated with such closed-form instructions exhibited limitations in real-world user scenarios, prompting the development of an alternative approach. To address these constraints, the concept of open-domain instruction tuning [30, 49] is conceived. OpenAI pioneered this approach by employing human annotators to compile a corpus of real-world question-answer datasets. These datasets form the foundation for training a reward model through reinforcement learning methodologies. The trained reward model then functions as a supervisory mechanism for further training instruction-oriented language models, such as InstructGPT [31] and Vicuna [57]. This innovation marks a significant advancement in the field, aiming to bridge the gap between LLM performance and real-world applicability by leveraging instruction data derived from authentic user interactions.
@@ 2.2 Large Vision Language Models
As the field of LLMs and instruction tuning undergoes rapid advancements, the academic research community is progressively focusing on integrating visual information into LLMs to facilitate visual instruction tuning. This emerging area of research has witnessed the development of various methodologies [22, 21, 5, 3, 54] based on foundational vision-language models [33, 47, 40, 19] and diverse LLM architectures [57, 43, 3]. In particular, LLaVA [22] pioneers the integration of an LLM with a CLIP vision encoder to construct a vision language model, demonstrating remarkable capabilities in image-text dialogue tasks through pretraining alignment strategies and targeted instruction tuning. Subsequent investigations have sought to refine visual instruction tuning by enhancing the quality and variety of the datasets used during the pre-training and fine-tuning phases. Building upon these advancements, recent studies like LLaVA-v1.5 [21] and ShareGPT4V [5] have achieved notable success in general vision-language comprehension, showcasing their ability to undertake complex question-answering tasks. This progression underscores the importance of sophisticated data handling and model tuning strategies in developing effective vision-language models.
@@ 3.1 Architecture Overview
The proposed SQ-LLaVA model (see Fig. 3) consists of four main components: 1) A pre-trained vision encoder CLIP-ViT [33] that extracts a sequence embedding of image tokens $Z_{v}$ for an input image $X_{v}$; 2) A prototype extractor $\phi(\cdot)$ learning visual clusters to enhance the original image tokens; 3) A trainable projection block $W(\cdot)$ with two linear layers to map the enhanced image tokens to the language domain tokens $H_{v}$, handling the dimension misalignment between the vision and language domain; and 4) Our LLM backbone $f(\cdot)$ implemented by the pre-trained Vicuna [57] to predict the next token upon the previous embedding sequence. Given the input question $X_{q}$ and answer $X_{a}$, a word embedding matrix is used to map them to contextual embeddings $H_{q}$ and $H_{a}$, and the distribution over $H_{a}^{(i+1)}$ can be obtained following the auto-regressive model as:
@@
where $\theta$ represents all the trainable parameters in our model, $\sigma(\cdot)$ is a softmax function, and $f(\cdot)$ outputs the last token embedding of the whole sequence. We denote $p_{\theta}$ as the prediction probability for the anticipated answer token $H_{a}^{(i+1)}$ at the position $i+1$, conditioning on the input image token embedding $H_{v}$, the question token embedding $H_{q}$, and the previous answer token embeddings $H_{a}^{(1:i)}$. As shown in Eq. (1), the proposed SQ-LLaVA applies the language model $f(\cdot)$ to model $p_{\theta}$ given by $H_{v}$, $H_{q}$, and $H_{a}^{(1:i)}$. Existing visual instruction tuning methods [21, 5] are only able to predict $H_{a}$, yet cannot fully exploit the rich semantic clues within $H_{q}$. In this study, we propose visual self-questioning instructions to guide LLMs in capturing the visual content within questions.
@@ 3.2 Visual Self-questioning Instruction
In broad real-world scenarios, proactively asking a question requires more understanding and background knowledge than answering [41]. Similarly, this work proposes visual self-questioning to encourage the LLM to discover deeper vision-language alignment and improve the overall instruction-following performance. Particularly, SQ-LLaVA treats questioning as a new learning objective beyond answering, which, to the best of our knowledge, is the first practice in the field of visual instruction tuning. While the current vision-language model can ask questions [1], such skill is still learned from question-answering through instruction tuning. However, our proposed method shows that the decoder-based LLM has the potential to learn more skills, such as how to ask questions spontaneously when a unique instruction token is given (e.g., we define [vusr] in our work). Furthermore, visual self-questioning can potentially improve general vision-language understanding. To be specific, as shown in Fig. 1, there are a certain amount of questions containing more meaningful image-related information than answers in the existing visual instruction data [21, 5]. Thus, we hypothesize that the vision-language understanding can be improved once the LLM learns how to predict relevant questions about a given image.
@@
$X_{c}^{j}=\left\{\begin{matrix}(\texttt{[usr]},X_{q}^{(j)},\texttt{[aswr]},X_{a}^{(j)})&~{}~{}j=1\mathrm{~{}or~{}}j>1,R<\delta\\ ~{}(\texttt{[{\color[rgb]{0.5625,0.66796875,0.86328125}\definecolor[named]{pgfstrokecolor}{rgb}{0.5625,0.66796875,0.86328125}vusr}]},X_{q}^{(j)},\texttt{[aswr]},X_{a}^{(j)})&j>1,R>\delta\end{matrix}\right.,$ (2)
@@
SQ-LLaVA performs zero-shot questioning without any in-context knowledge or human language instruction since the only instructional prompt is a unique token [vusr]. After visual self-questioning on instruction data with various question formats $X_{q}$, the questions sampled by SQ-LLaVA are more diversified than GPT4-V (as shown in Fig. 5), since the LLM has learned alignment between image and questions, which is different from previous works [1, 48]. Specifically, previous general-purpose vision language models such as GPT4-V [1] can generate questions based on a given image, but it requires explicit language instruction such as “Ask complex questions that are relevant to the content in the image”. Accordingly, the quality of generated questions highly relies on prompt engineering. Also, self-instruct [48] utilizes in-context learning to prompt the LLM to ask specific questions.